OpenStudy (anonymous):

what is 3000 in scientific notation

OpenStudy (anonymous):

how many zeros do you count?

OpenStudy (anonymous):

\[\large 3\times 10^{\text{number of zeros}}\]

OpenStudy (skullpatrol):

$$10^1=10$$ $$10^2=10*10=100$$ $$10^3=10*10*10=1,000$$

OpenStudy (skullpatrol):

$$3,000=3*1,000$$

OpenStudy (skullpatrol):

https://www.youtube.com/watch?v=i6lfVUp5RW8